# From shape to fate: making bacterial swarming expansion predictable

### v1

@@ 3 Discussion
Swarming colonies are often cited as emblematic examples of emergent collective behaviour, yet most analyses remain anchored in static descriptions of highly structured fronts. Here we recast the same imaging data into a predictive formulation of swarming dynamics. Rather than treating expansion as a sequence of images, the colony edge is modelled as a dynamical object evolving within a low–dimensional morphological state space. By assembling the SwarmEvo dataset, introducing TexPol–Net for boundary measurement, and developing Morpher as a morphology–level dynamical model with structural memory, the expansion of Enterobacter sp. SM3 becomes a forecasting problem defined at the level of geometry, rather than a descriptive account of shape.
This reframing rests on an explicit separation between measurement, state construction and dynamics. Time–lapse videos are first reduced to sequences of boundary–resolved masks, defining the measured observables of the system. TexPol–Net operates at this stage not simply as a segmentation network, but as a boundary measurement operator that stabilizes the morphological degrees of freedom of the colony front. By suppressing diffuse edges, preserving thin protrusions and decoupling the front from background artefacts, TexPol–Net defines a consistent coordinate system in which small geometric variations retain temporal coherence. Improvements in boundary accuracy therefore propagate nonlinearly into forecasting performance, as every measurement error reshapes the subsequent trajectory in state space. In this sense, TexPol–Net establishes the state variables on which dynamics can be meaningfully learned, rather than merely refining per–frame masks.
This sensitivity becomes most apparent when segmentation backbones are required to resolve near–equivalent masks. On SwarmEvo, YOLOv12 differs from TexPol–Net by less than one mAP50:95 point (91.81 vs. 92.48; Fig. 3c), yet this minor discrepancy at the measurement stage expands once the masks define the state variables for autoregressive forecasting. When Morpher is held fixed, boundary deviations in YOLOv12 inputs, including slight edge displacement and reduced high–curvature detail, propagate through successive rollouts, producing gradual front drift and a measurable loss of curvature that reaches a 2.4–3.1 IoU gap after 20% prediction (Fig. 2.3c). TexPol–Net masks, by contrast, sustain consistent front localization and curvature over extended horizons, limiting geometric degradation during temporal propagation (Fig. 3b). SAM/SAM2 accentuate this contrast by suppressing distal structure and biasing contours toward the colony core, effects that are already evident at the segmentation stage (Fig. 3). These comparisons indicate that once the colony edge is treated as the dynamical state, boundary fidelity directly governs whether temporal prediction remains stable or undergoes progressive error growth.
Morpher operates on this measurement–defined state space to model the evolution of the front. Unlike generic video prediction architectures, Morpher is constrained to propagate contours and to encode dynamics in terms of geometric quantities such as curvature and radial displacement (Fig. 2.5). This design choice places Morpher explicitly in the role of a dynamical model, rather than an image generator. Benchmarking against state–of–the–art sequence models shows that enforcing this geometric alignment improves both region–level overlap and boundary fidelity, while preserving anisotropic branching despite the absence of texture information. Ablations across backbones, inference strategies and the Morphon module reinforce a mechanistic interpretation of the model behaviour (Fig. 2.5b). Autoregressive decoding consistently yields more realistic front propagation than parallel one–shot prediction, reflecting the incremental nature of physical growth. Architectures equipped with attention and persistent structural memory, particularly the Transformer with Morphon, best preserve fingering and radial symmetry breaking, indicating that long–range couplings are integral to swarming dynamics.
Varying the observation ratio provides further insight into the stability of morphology–level forecasting (Fig. 2.5c). As the observed fraction increases from 50% to 90%, coarse geometric metrics such as mIoU, HD–based distances and ASSD improve and then saturate. Once slightly more than half of the trajectory is observed, additional frames contribute little to recovering the global outline or average advance speed. In contrast, temporal and angular descriptors reveal where prediction remains sensitive. TCI exhibits measurable dependence on the forecasting window, indicating that small phase shifts in accelerations and pauses perturb fluctuation profiles even when boundaries remain well aligned.
Angular metrics expose a more stringent regime. The non–monotonic dependence of $|\Delta\mathrm{NAS}|$ on observation ratio suggests that anisotropy errors are governed less by the amount of historical data than by how the prediction window intersects late–stage sectoring and branch rearrangements. In this regime, minor mismatches in timing or amplitude reshape the angular power spectrum even when global geometry is preserved. The behaviour of $|\Delta\mathrm{H}_{2}|$ sharpens this interpretation. Isolating the second harmonic renders the metric sensitive to the redistribution of curvature energy across Fourier modes. At high observation ratios, prediction targets are dominated by fine protrusions and tip instabilities, where small deviations in low–frequency content amplify into substantial harmonic differences. These results indicate that anisotropic fingering and harmonic organization represent the most demanding aspects of swarming to forecast, and they suggest a fundamental sensitivity of morphology–only dynamics to small temporal errors.
Several limitations delineate avenues for extension. SwarmEvo focuses on a single strain, Enterobacter sp. SM3. Although its behaviour spans multiple regimes, swarming encompasses a broad spectrum of species and substrates whose statistics may differ substantially. The present representation further collapses colonies into quasi–two–dimensional fronts. While such projections preserve the dominant geometric degrees of freedom and are standard in swarming studies, height and density variations enter only implicitly. Morpher is trained without explicit biophysical constraints; the absence of hand–crafted rules affords flexibility, yet extrapolation into sparsely sampled regimes may benefit from mechanistic priors. Finally, forecasting is treated as an open–loop problem. Closed–loop experiments in which predictions guide perturbations would provide a natural next step toward control.
Beyond swarming, this work establishes a general methodology for forecasting dynamics in systems where appearance is uninformative but geometry is structured. By formalizing expansion as a measurement–defined dynamical system—where boundary measurement defines the state space and forecasting reduces to trajectory evolution rather than image prediction—the framework clarifies how segmentation, state construction and temporal inference interact. Representing videos as trajectories in a geometry-aligned morphological state space enables principled comparison of sequence models, separates measurement error from dynamical uncertainty, and supports interpretable metrics linked directly to branching, anisotropy and harmonic organization. These principles extend to other forms of collective motility, biofilm spreading and tissue invasion, where dynamics unfold through shape rather than texture. By enabling long–horizon forecasts and identifying features that are intrinsically sensitive to timing, this framework transforms colony expansion into a quantitative, methodologically grounded basis for intervention, providing a foundation for predictive modelling and control of microbial collectives in complex environments.
@@ 4 Methods
4.1 Experimental workflow, imaging, and mask generation
@@
The experimental workflow followed Fig. 1. A single colony of Enterobacter sp. SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and incubated at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm for approximately 16 h to obtain a high-density culture. A $5\text{--}8\,\mu\mathrm{L}$ aliquot of the overnight culture was spotted onto the center of a freshly prepared swarming agar plate (LB medium supplemented with 0.5–0.8% agar), poured to a uniform thickness of 3–4 mm to promote consistent hydration and nutrient diffusion across plates. The inoculated plates were incubated at $30\,^{\circ}\mathrm{C}$ and approximately 90% relative humidity for 4–6 h to activate swarming motility, and then transferred to a time-lapse imaging chamber maintained between $28\text{--}37\,^{\circ}\mathrm{C}$ and 80–92% relative humidity. Humidity was kept below the condensation threshold to prevent droplet formation on the agar surface. Surface stiffness and inoculum volume were varied across experiments to introduce controlled physical differences and generate diverse colony expansion conditions.
@@
Imaging was performed using a vertically mounted high-resolution digital camera with uniform LED illumination to minimize specular reflection and shadowing artifacts. Frames were recorded at one-minute intervals throughout colony expansion and continued until the colony front reached the plate boundary or no further measurable growth was observed. Sequences were stored at native spatial resolution with associated timestamps.
For downstream analysis, each sequence was segmented using TexPol–Net to obtain pixel-level colony masks. These mask sequences formed a morphology-consistent representation used for training and evaluating the forecasting models. Training and validation datasets were assembled from non-overlapping experimental runs collected under distinct growth conditions to assess generalization across colony expansion regimes.
4.2 Recovering front texture and geometry
Swarming colony images present two coupled challenges for convolutional backbones. At the micro-structural level, boundaries are uncertain and textures vary across scales, so high-frequency evidence is easily attenuated when features are aggregated over local receptive fields. At the meso- to macro-structural level, growth follows a disc-like geometry with coherent radial propagation, producing long-range dependencies that are poorly captured by standard kernels. Addressing both phenomena with a single mechanism tends to inflate complexity or compromise either edge fidelity or global organization.
The proposed design decouples the challenges of texture variation and geometric organization while maintaining a unified attention framework. The Texture-Edge Attention (TEA) module, illustrated in Figure 6a, is devised to enhance the stability and clarity of fine-scale features that are often blurred by conventional filters. It integrates three cooperative components: a local depthwise branch that preserves intra-channel spatial detail, a multi-dilated branch that captures textures over multiple scales, and an edge-sensitive branch initialized with a discrete Laplacian kernel to introduce a high-pass prior at the beginning of training. Channel-wise and spatial attention mechanisms further purify the fused representation by reducing redundancy and highlighting informative regions, while a conditional residual pathway ensures optimization stability when feature dimensions match. Detailed formulations and implementation details for both modules are presented in Section S2.
The Polar–Context Attention (PCA) module, illustrated in Figure 6b, complements TEA by embedding a radial geometric prior directly into the feature representation. The module first compresses feature channels to eliminate redundancy and then performs a polar transformation that maps features into $(\rho,\theta)$ space, where depthwise dilated convolutions operate along radial and angular directions consistent with the colony’s outward propagation. A large-kernel Cartesian branch provides broad contextual coverage, while a local branch preserves fine structural fidelity. Channel- and spatial-attention mechanisms further refine the fused representation, enhancing the coherence between local detail and global geometry. A conditional residual connection stabilizes training when feature dimensions coincide.
TEA and PCA provide complementary inductive biases: TEA sharpens edges and preserves high-frequency texture amid blurred, irregular boundaries, whereas PCA captures geometry-aligned, long-range dependencies characteristic of near–concentric–ring radial growth. Operating both modules at a compact internal width with shared normalization and activation yields a representation that is simultaneously edge-aware, scale-robust, and geometry-aligned, without incurring prohibitive computational cost.
4.3 TexPol–Net
Texture- and geometry-aware instance segmentation. To recover accurate colony fronts from swarming imagery, we developed the Texture-Polar Network (TexPol–Net), an instance segmentation architecture that integrates fine-scale texture encoding with geometry-aligned contextual reasoning. This design addresses two coupled challenges in swarming colonies: diffuse and low-contrast boundaries at the local scale, and near–concentric radial organization at the global scale. TexPol–Net draws inspiration from efficient one-stage detectors such as YOLO [41, 35] and the prototype-based mask formulation of YOLACT [3], while introducing texture- and geometry-aware mechanisms tailored to biological front segmentation (Fig. 3a).
Hierarchical backbone with complementary attentions. TexPol–Net employs a five-stage hierarchical convolutional backbone with progressive downsampling. At each stage, Texture-Edge Attention (TEA) and Polar–Context Attention (PCA) are jointly embedded to serve complementary roles. TEA operates predominantly on shallower representations to preserve boundary sharpness and local texture variation, whereas PCA increasingly emphasizes geometry-aligned contextual aggregation in deeper layers, enabling the capture of long-range radial propagation patterns.
Bidirectional multi-scale feature fusion. The neck adopts a PANet-style bidirectional fusion strategy [25] to integrate semantic and spatial information across scales. Top–down and bottom–up pathways are interleaved with PCA modules, ensuring that polar consistency is preserved during resolution recovery. A final TEA module is applied prior to prediction to recalibrate channel responses and reinforce fine-scale texture cues.
Dense prediction and prototype-based mask assembly. The prediction head employs fully convolutional branches on feature levels $P3$, $P4$, and $P5$ to predict class scores, bounding boxes, and instance-specific mask coefficients. Instance masks are generated using a YOLACT-style prototype formulation: a lightweight Protonet produces k=32k{=}32 shared prototypes, and each detection predicts a coefficient vector for linear combination. After non-maximum suppression, masks are generated, cropped, and thresholded within the retained bounding regions.
Loss. During training, supervision is applied to all prediction heads in each mini-batch. The overall objective is defined as a weighted linear combination of the bounding box regression, classification, distribution focal loss (DFL), and mask segmentation terms:
$\mathcal{L}_{\mathrm{seg}}=\lambda_{\mathrm{box}}\,\mathcal{L}_{\mathrm{box}}+\lambda_{\mathrm{cls}}\,\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{dfl}}\,\mathcal{L}_{\mathrm{dfl}}+\lambda_{\mathrm{mask}}\,\mathcal{L}_{\mathrm{mask}}.$
The box loss penalizes geometric misalignment between predicted and ground-truth boxes, the classification loss supervises category confidence, and DFL refines localization by regularizing the underlying discrete distribution. The segmentation loss enforces consistency between predicted and ground-truth instance masks. All loss terms are combined linearly with fixed weights, which are kept constant across experiments.
4.4 Morpher
Mask-space forecasting. To model the temporal evolution of swarming colony morphology, we developed Morpher, a forecasting network that operates on segmentation masks rather than raw image intensities. This choice follows the biology of swarming expansion, where growth is expressed through the geometry and advancement of the colony front. Mask-level representation isolates occupied regions and evolving boundaries, focusing the model on front curvature and large-scale expansion while reducing sensitivity to illumination and imaging variability (Fig. 2.5).
Spatial encoder and latent sequence. Given an input mask sequence, a multi-scale encoder extracts a hierarchy of feature maps $\{f_{1}^{(t)},f_{2}^{(t)},f_{3}^{(t)},f_{4}^{(t)},f_{5}^{(t)}\}$ at each time step via progressive downsampling $(H\!\to\!H/2\!\to\!H/4\!\to\!H/8\!\to\!H/16\!\to\!H/32)$. In parallel, each frame is compressed into a latent descriptor $z_{t}\in\mathbb{R}^{256}$, yielding a compact temporal representation for forecasting. For temporal modeling, the observed segment is represented as a latent tensor of shape $(B,T_{\mathrm{obs}},256)$, where the latent dimensionality is fixed by the shared spatial encoder. Sinusoidal time encodings are added to the latent sequence prior to the temporal module to provide a uniform representation of temporal order.
@@ 4.4 Morpher
Autoregressive temporal evolution. Morpher forecasts in an autoregressive loop. After observing $T_{\mathrm{obs}}$ frames, the model predicts $\hat{z}_{T_{\mathrm{obs}}+1}$, decodes it into a mask, and re-encodes the prediction as the next input. This feedback couples future states to the geometry produced at the previous step and helps maintain boundary continuity over long horizons. Forecasting is evaluated under two inference regimes in matched settings. In the autoregressive regime, predictions are generated stepwise over the horizon, with each predicted mask fed back as the next input. In the parallel regime, the model predicts the full forecasting horizon conditioned on the observed segment and the terminal latent state.

### v2

@@ 3 Discussion
Swarming is a striking example of emergent collective behavior [66, 25], yet its advancing front remains treated as a descriptive pattern rather than a predictive dynamical state. Here we recast Enterobacter sp. SM3 expansion as a predictive problem defined at the colony edge. This shift is biologically consequential: in inflamed intestinal environments, the advancing front is the interface through which swarming reshapes oxygen availability, niche structure, and downstream community organization (Figure 1[14]. The question is therefore not simply how a colony looks, but whether the future position and organization of its front can be inferred from its present state.
A central result is that this future is not cleanly indexed by nominal assay conditions. Trajectories remain broadly overlapping across temperature, humidity, and agar groupings (Figure 2a,b), and within- and across-condition distances are only weakly separated (Figure 2c,d). Prediction therefore depends on colony-specific morphology rather than on coarse labels. This helps explain why appearance-based extrapolation fails: the relevant information is carried by front geometry, protrusions, and temporal continuity, not by condition identity or framewise visual similarity [65].
That observation makes measurement decisive. TexPol–Net matters not simply because it improves segmentation, but because it defines the state on which dynamics are learned. Once the colony edge is treated as the evolving variable, small boundary errors are no longer cosmetic (Figure 3c–e). A sub-point difference in segmentation quality between TexPol–Net and YOLOv12 (92.48% vs. 91.81% mAP50:95) expands into a 2.4–3.1 IoU loss after autoregressive rollout (Figure 3c and Figure 4c). Stable forecasting therefore requires boundary fidelity at the measurement stage.
Within that state space, Morpher behaves, as schematized in Figure 5a, as a front-dynamics model rather than an image predictor. Its advantage over generic video architectures lies not only in overlap, but in preserving front localization and anisotropic branching over long horizons (Figure 4a,b). The temporal ablations sharpen this point. Autoregressive decoding consistently outperforms parallel prediction, indicating that swarming is better modeled as incremental state propagation than as one-shot future synthesis (Figure 5b). More importantly, the model that minimizes propagation error is not the one that best preserves morphology: GRU achieves the lowest RMSE, whereas Transformer with Morphon best maintains boundary fidelity and anisotropic structure (Figure 5b). The harder part of swarming prediction is therefore not coarse advance alone, but retention of branch history and late-stage front organization [67].
The observation-ratio and leave-one-out analyses place that difficulty more precisely. Global geometric metrics improve and then saturate with longer observation, whereas angular descriptors remain sensitive to late rearrangements of the front (Figure 5c). At the same time, leave-one-out evaluation shows that the model preserves not only mask overlap but also the temporal organization of effective radius, perimeter, and propagation velocity across held-out trajectories (Figure 6b,d–h). Together, these results indicate that the dominant residual challenge lies in anisotropic branching and phase-sensitive reorganization, not in recovery of coarse colony extent.
Several limitations define the next steps. While leave-one-out evaluation already indicates that the model captures transferable dynamics under limited data, SwarmEvo is currently restricted to a single strain, Enterobacter sp. SM3, with broader taxonomic and substrate diversity remaining as a natural extension to assess the generality of these dynamics. The present representation is quasi-two-dimensional, so height and density enter only indirectly. Morpher also operates without explicit biophysical constraints; incorporating mechanistic priors may further improve extrapolation in sparsely sampled regimes. Finally, forecasting remains open-loop. Closed-loop perturbation experiments guided by model predictions would provide the most direct test of whether predictive morphology can be translated into controllable intervention at the swarming front.
More broadly, this framework suggests a route for living systems in which shape carries the essential dynamics [19]. By separating measurement, state construction, and temporal evolution, it becomes possible to ask which aspects of collective behavior are predictable, which remain unstable, and which are most sensitive to perturbation. For swarming, this turns morphology from a descriptive readout into a dynamical variable that can be compared, forecasted, and ultimately acted upon.
4 Conclusions
Boundary-resolved morphology provides a workable state space for forecasting Enterobacter sp. SM3 swarming. SwarmEvo supplies the time-lapse trajectories, TexPol–Net recovers fronts with sufficient fidelity to define the predictive state, and Morpher advances those states through autoregressive rollout with structural memory. Under the 80% observation / 20% prediction protocol, Morpher achieves the best overall performance (mIoU 95.42%, HD95 10.61 px, ASSD 3.93 px), outperforming the strongest baseline by 5.4% in overlap and reducing HD95 and ASSD by 42.0% and 55.7%, respectively. Just as importantly, the study shows why forecasting is feasible only in a geometry-first formulation: once morphology is measured accurately, swarming expansion can be treated as a predictable dynamical process rather than an appearance-driven one. This establishes a basis for quantitative comparison of swarming trajectories across environments and for future predictive control of advancing microbial fronts.
@@ 5 Methods
5.1 Swarming assay, imaging, and dataset construction
@@
The experimental workflow followed Figure 1. A single colony of Enterobacter sp. SM3 was transferred from an LB agar plate into LB broth (10 g/L tryptone, 5 g/L yeast extract, and 5 g/L NaCl) and cultured overnight at $37\,^{\circ}\mathrm{C}$ with shaking at 200 rpm. A $5\text{--}8\,\mu\mathrm{L}$ aliquot was inoculated at the center of a freshly prepared swarming plate (LB with 0.5%, 0.6%, or 0.7% agar; plate thickness 3–4 mm), incubated at $30\,^{\circ}\mathrm{C}$ and $\sim 90\%$ relative humidity for 4–6 h to activate swarming, and then transferred to a time-lapse imaging chamber maintained at $27\,^{\circ}\mathrm{C}$, $30\,^{\circ}\mathrm{C}$, or $33\,^{\circ}\mathrm{C}$ and $86\%$, $90\%$, or $94\%$ relative humidity. These temperature, humidity, and agar conditions span the permissive regime of SM3 swarming [47, 11, 27] and were varied across discrete levels to probe model generalization within physiologically relevant growth states. Humidity was kept below the condensation threshold to avoid droplet formation on the agar surface. Experimental conditions were varied to generate distinct expansion regimes.
@@
Images were acquired with a vertically mounted high-resolution digital camera under uniform LED illumination. Frames were recorded every minute until the colony reached the plate boundary or no further measurable expansion was observed. The resulting Swarming Morphogenesis Evolution (SwarmEvo) dataset comprised 1,971 annotated images for segmentation and 276 long time series for temporal modeling. All recordings were stored at native resolution ($1250\times 1250$ px) with timestamps.
For downstream analysis, each sequence was converted into boundary-resolved colony masks using TexPol–Net. These masks defined the state space for forecasting. Training and validation partitions were split at the sequence level, with no colony contributing trajectories to both partitions. Exact dataset composition and augmentation procedures are provided in the Supplementary Information.
5.2 TexPol–Net for boundary-resolved state construction
TexPol–Net was designed to recover colony fronts under two coupled conditions: diffuse, low-contrast boundaries at the local scale and near-concentric radial organization at the colony scale. The network therefore combines two complementary modules. Texture–Edge Attention (TEA) preserves fine boundary texture through local depthwise filtering, multi-scale dilated texture encoding, and an edge-sensitive high-pass prior. Polar–Context Attention (PCA) embeds a geometry-aligned prior by combining local features, large-kernel Cartesian context, and a polar branch operating in $(\rho,\theta)$ coordinates. Full module formulations and implementation details are provided in the Supplementary Information.
As shown in Figure 3a, the full architecture follows a prototype-based one-stage instance segmentation design [52, 5]. A five-stage hierarchical convolutional backbone progressively downsamples the input while interleaving TEA and PCA blocks, allowing fine boundary cues and geometry-aligned context to be encoded jointly. The resulting multi-scale features are fused by a PANet-style bidirectional neck [36, 38], in which PCA is retained during top-down and bottom-up aggregation to preserve polar consistency across resolutions. Prediction is performed at feature levels $P3$, $P4$, and $P5$ through dense heads for class scores, bounding boxes, and instance-specific mask coefficients. In parallel, a lightweight Protonet produces $k=32$ shared prototypes, which are linearly combined with the predicted coefficients and then cropped and thresholded to produce final colony masks.
Training uses a composite segmentation objective,
$\mathcal{L}_{\mathrm{seg}}=\lambda_{b}\mathcal{L}_{b}+\lambda_{c}\mathcal{L}_{c}+\lambda_{d}\mathcal{L}_{d}+\lambda_{m}\mathcal{L}_{m}.$
Here $\mathcal{L}_{b}$, $\mathcal{L}_{c}$, $\mathcal{L}_{d}$, and $\mathcal{L}_{m}$ denote box, classification, distribution focal, and mask losses, respectively. All weights were fixed across experiments.
5.3 Morpher for morphology forecasting
Morpher forecasts swarming expansion in mask space rather than image space. This formulation follows the biology of the problem: what propagates is the front, not visual appearance. Each input mask sequence is encoded by a shared multi-scale spatial encoder into framewise latent descriptors $z_{t}\in\mathbb{R}^{256}$, together with intermediate feature maps used later for decoding. Sinusoidal temporal encodings are added before temporal modeling.
As summarized in Figure 5a, Morpher consists of four coupled components: a multi-scale spatial encoder, a temporal sequence model, a Morphon memory block, and a multi-scale decoder. Observed masks are first compressed into a compact latent sequence, while encoder-side intermediate features are retained for later reconstruction. Future evolution is then predicted in latent space. In the autoregressive setting, each decoded prediction is re-encoded and fed back as the next input, so that the forecast proceeds through stepwise updates rather than one-shot generation. Morphon operates on the observed latent history by cross-attention with a learnable query derived from the aggregated observation state, and injects the retrieved structural memory through a learnable gate $\alpha\in(0,1)$. A multi-scale decoder finally reconstructs the predicted mask while reinjecting encoder features to preserve peripheral protrusions and fine curvature.
@@ 5.3 Morpher for morphology forecasting
Forecasting is performed autoregressively unless otherwise stated. After observing $T_{\mathrm{obs}}$ frames, the model predicts the next latent state, decodes it into a mask, re-encodes that prediction, and feeds it back for subsequent steps. This stepwise rollout couples future predictions to the geometry produced at the previous step and stabilizes boundary continuity over long horizons. Parallel prediction is used only in matched ablations.